OpenStudy (anonymous):

g(x) = 2x2 + 7 h(x) = 4x2 + 7 Find: (g - h)(5) A= -43 B.=-57 C=-50 D=164

OpenStudy (anonymous):

@johnweldon1993 your literally the only one who ever helps me.... Any idea?

OpenStudy (sidsiddhartha):

hint (g-h)(5)=g(5)-h(5)

OpenStudy (anonymous):

what do i need to do with this hint?

OpenStudy (sidsiddhartha):

what is the value of g(5) its just =2*5^2+7=57 and h(5)=4(5^2)+7=107 so g(5)-f(5)=57-107=-50

OpenStudy (anonymous):

ohhh okay... Thanks when it comes to math im like a child.